HUBBLE SPACE TELESCOPE DAILY REPORT #2099 PERIOD COVERED: 0000Z (UTC) 04/09/98 - 0000Z (UTC) 04/10/98 Daily Status Report as of 100/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled with no anomalies. 1.3 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.5 Completed FGS 7485 (Dynamical Masses for the Stars in the Pre-Main-Sequence Spectroscopic Binary 045251+3016) The Fine Guidance Sensors were used to make astrometric observations of the pre-main-sequence, spectroscopic, binary STAR-04560 in order to obtain its dynamic mass. An excellent single-lined spectroscopic orbit is already completed, which provides all of the essential orbital elements except inclination angle and mass ratio. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.7 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.8 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed FGS 7492 (The Distances to Dwarf Novae, and the Calibration of the Technique of Infrared Spectroscopic Parallax) The Fine Guidance Sensors was used to make astrometric observations of the dwarf nova U-Gem and several other reference stars. The distances to dwarf novae are poorly known. The distance to the object can be measured using a technique of infrared spectroscopic parallax. The observations were completed as scheduled with no anomalies. 1.10 Completed Two Sets of NIC/1/2/3 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of Super Nova SN1998E in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Three Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.12 Completed Three Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed WF/PC-2 7627 (WF/PC-2 Cycle 7 Astrometric Monitor) The WF/PC-2 was used to make observations of the large globular Cluster Omega Cen as part of a Cycle 7 astrometric monitor and calibration of the CCDs. This program will verify accuracy and stability of geometric transformation and relative astrometry solution. The observations were executed as scheduled and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: Using ROP SR-1, the EDAC counter was cleared at 099/1838Z. ESTR-2R reconditioning was completed at 100/0655Z as directed by ROP RD-7. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH